Final problem question for me
(-8)(-5) + 7^ {2)

someone please help me

Answer by checkley77(12844)

(-8)(-5) + 7^ {2)
-8 TIMES -5=40 [NOTE: 2 NEGATIVES MULTIPLIED IS A PLUS]
7^2=7*7=49
THUS:
40+49=89 ans.